OpenStudy (chosenmatt):

what is -3 2/3 times (-2 1/4)

OpenStudy (matlee):

-11/3 * -9/4= 99/12= 9 1/4

OpenStudy (anonymous):

I'll tell you. Give me a medal afterwords.

OpenStudy (anonymous):

- 99/12

OpenStudy (matlee):

its posotive

OpenStudy (anonymous):

oh sorry.